OpenStudy (anonymous):

Which is a rational number? A: sqrt 10 B: sqrt 24 C sqrt 36 D: sqrt 44

OpenStudy (anonymous):

which number isn't a decimal when you find the square root

OpenStudy (anonymous):

36?

OpenStudy (anonymous):

yup so it's rational

OpenStudy (anonymous):

Oh XD Ok thanks!!

OpenStudy (anonymous):

Your welcome : )

OpenStudy (solomonzelman):

\[1=(±1)^2\]\[4=(±2)^2\]\[9=(±3)^2\]\[16=(±4)^2\]\[25=(±5)^2\]\[36=(±6)^2\] should \(\Large\color{blue}{ \bf remember }\) these perfect squares up to 16 at least.